OpenStudy (anonymous):

x''=(x')^2 how can you reduce to a first order diff eq?

OpenStudy (perl):

x'' (t) - x' ^2 = 0

OpenStudy (anonymous):

so is that separable to y=-1/(x+C1) ?

OpenStudy (perl):

not sure, one sec

OpenStudy (perl):

let x'(t) = v(t)

OpenStudy (perl):

let start over x''(t) = x'(t)^2

OpenStudy (perl):

if x'(t) = v(t), then x''(t) = v'(t)

OpenStudy (perl):

now we have v'(t) = v(t)^2, which is first order

OpenStudy (anonymous):

right so it would be v'(t)+v(t)^2=0

OpenStudy (perl):

minus

OpenStudy (perl):

how about we divide both sides by v(t)^2

OpenStudy (perl):